OpenStudy (anonymous):

find the limit as x approaches 1 for (x^2 - x^(1/2)) / (1-x^(1/2))

OpenStudy (anonymous):

\[(x ^{2}-\sqrt{x})/(1-\sqrt{x})\]

OpenStudy (experimentx):

\[ x^2 = (\sqrt x)^4 \] take sqrt x common, expand using a^3-b^3, do cancelling operation and put values to get limit

OpenStudy (anonymous):

ahh okay that makes sense, thank you !!